OpenStudy (anonymous):

Please help me find the integral of ∫(12/∛(t^2 )-5/e^2t - 7/t^2 )dt

OpenStudy (anonymous):

The original equation

OpenStudy (anonymous):

Do you know that you can split up the integral?

OpenStudy (anonymous):

oops sorry, the first one should have a 12 in front of the integral, I misread the equation above, besides that it's correct though.

OpenStudy (anonymous):

I will rewrite it to avoid confusion

OpenStudy (anonymous):

The first integral is 12t^-2/3

OpenStudy (anonymous):

\[ \large 12\int t^{-\frac{2}{3}} dt -5 \int e^{-2t}dt -7 \int t^{-2} dt \]

OpenStudy (anonymous):

correct @nishika17 , my eyes are getting worse it seems (-;

OpenStudy (anonymous):

Ya now it can be solved by simple integral ormula

OpenStudy (anonymous):

the middle integral requires a substitution, or an advanced guess.

OpenStudy (anonymous):

\[\int\limits e^-2t.dt put -2t= u differentiating on both sides we get -2 .dt= du or dt= -du/2\]

OpenStudy (anonymous):

\[\int\limits -1/2 e^u. du = -1/2 e^u or -1/2e^-2t\]

OpenStudy (anonymous):

\[thats the answer of \middle part o \int\limits\]

OpenStudy (anonymous):

Thank you all so much for you great help!

OpenStudy (anonymous):

welcome